OpenStudy (anonymous):

User: Assume 3 in every 3000 students at the local community college have to quit due to serious health issues. An insurance company offers them ma$12,000 policy for $50 a year. What is the amount the insurance company should expect to make on average on every student that pays?

OpenStudy (anonymous):

Each year the company expects to pay out (3)(12000)= $36000 It expects to collect (50)(3000) = $150,000 per year the difference is $114000 per year.

OpenStudy (anonymous):

Thanks, they said the answer is wrong.

OpenStudy (kropot72):

The probability that a student claims is 3/3000 = 1/1000 The average loss per student is \[12000\times\frac{1}{1000}=12\ dollars\] The income per student is $50. Therefore the average profit per participating student is 50 - 12 = $38

OpenStudy (anonymous):

Did you divide the $114,000 gain by the number of students to get a per student estimate? That gives $38.

OpenStudy (kropot72):

@douglaswinslowcooper Your calculation assumes that every student participates. Although your method can arrive at a correct answer, I believe that mine is more in keeping with the question: "What is the amount the insurance company should expect to make on average on every student that pays?"